OpenStudy (anonymous):

Find (dy/dx) when y=10x^5 * sin^3 of x^3 * cos^4 of (3x+7) @Calculus1

OpenStudy (anonymous):

can I drink you ?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

you are soda right ?

OpenStudy (jamesj):

The 'trick' of course is to think of this function as a product of several functions. Let u = u(x) = 10x^5 v = v(x) = sin^3(x^3) w = w(x) = cos^4(3x+7) Then the original function is y = uvw and therefore the derivative y' = u'vw + uv'w + uvw' --- (*) Hence to calculate y' 1. Calculate each of the derivative u', v' and w'. You may need the chain rule for some of them 2. Then write down the expression (*) using those derivatives